# Assessing Open-world Forgetting in Generative Image Model Customization

### v1

@@ Abstract
Recent advances in diffusion models have significantly enhanced image generation capabilities. However, customizing these models with new classes often leads to unintended consequences that compromise their reliability. We introduce the concept of open-world forgetting to emphasize the vast scope of these unintended alterations, contrasting it with the well-studied closed-world forgetting, which is measurable by evaluating performance on a limited set of classes or skills. Our research presents the first comprehensive investigation into open-world forgetting in diffusion models, focusing on semantic and appearance drift of representations. We utilize zero-shot classification to analyze semantic drift, revealing that even minor model adaptations lead to unpredictable shifts affecting areas far beyond newly introduced concepts, with dramatic drops in zero-shot classification of up to 60%. Additionally, we observe significant changes in texture and color of generated content when analyzing appearance drift. To address these issues, we propose a mitigation strategy based on functional regularization, designed to preserve original capabilities while accommodating new concepts. Our study aims to raise awareness of unintended changes due to model customization and advocates for the analysis of open-world forgetting in future research on model customization and finetuning methods. Furthermore, we provide insights for developing more robust adaptation methodologies.
@@ 1 Introduction
Recent advancements in image generation have led to the development of remarkably powerful foundational models capable of synthesizing highly realistic and diverse visual content. Techniques such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), and more recently autoregressive models (Yu et al., 2022), Rectified Flows (Liu et al., 2023), and Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), have each contributed to significant progress in the field. These methods offer unique strengths in sample quality, diversity, and controllability. Among them, diffusion models have gained particular prominence due to their recent successes and growing influence, especially in enabling text-based image generation (Shonenkov et al., 2023; Ramesh et al., 2022) and complementary multimodal conditioning (Zhang & Agrawala, 2023; Mou et al., 2023), making them a key focus in current research and applications.
@@
Given the high-quality image generation capabilities of these models, a major focus of research has been on how to efficiently incorporate new content and adapt them to new tasks and domains. To address this challenge, various state-of-the-art transfer learning methods have been introduced. These include finetuning approaches such as DreamBooth (Ruiz et al., 2023b) and CustomDiffusion (Kumari et al., 2023), which allow models to learn new concepts effectively. Conditioning-based methods like ControlNet (Zhang & Agrawala, 2023) and IP-Adapters (Ye et al., 2023) enable precise control over generated images by incorporating additional guidance signals. Prompt methods like Prompt-to-Prompt (Hertz et al., 2023) and Textual Inversion (Gal et al., 2023) enable semantic image editing and learning new concepts without modifying the base model. Parameter-efficient techniques such as Low-Rank Adaptations (LoRA) (Hu et al., 2022) have shown great promise, allowing for rapid adaptation with minimal computational overhead (Blattmann et al., 2023; Shi et al., 2023b). These techniques enable models to learn new concepts effectively, even with few examples.
@@
These methods for adapting diffusion models (Ruiz et al., 2023b; Kumari et al., 2023) mainly rely on transfer learning and primarily focus on finetuning model weights to accommodate newly introduced data. However, they lead to unforeseen changes in the model’s behavior, which can have significant implications, altering the model’s existing knowledge, and skills, or the alignment between language and visual content within the network. The field of continual learning has long studied the issue of catastrophic forgetting in neural networks when these aim to adapt to new data (often referred to as new tasks) (Kirkpatrick et al., 2017; De Lange et al., 2021). Traditionally, it has focused on what we term closed-world forgetting, where evaluation is limited to a fixed set of classes encountered in previously learned tasks or skills (similar to concept editing). In contrast, modern foundation models introduce a new challenge: measuring open-world forgetting. Here, the model’s prior knowledge is vast, making it difficult to assess what has been forgotten or altered during the adaptation process.
@@
In this paper, we focus on two popular personalization methods, namely Dreambooth (Ruiz et al., 2023b) and CustomDiffusion (Kumari et al., 2023), for a case study of open-world forgetting. These techniques are especially relevant, as they only add very little new knowledge to the network: a single new concept represented by a small set of typically 3-5 images. Although one might expect that finetuning the model with such limited data would have minimal impact on the vast knowledge of the foundation model (e.g., Stable Diffusion), our analysis reveals that even these small updates can lead to highly detrimental consequences. As Figure 1 illustrates, finetuning can drastically alter the image representation of concepts seemingly unrelated to the training images. The complexity of the forgetting underscores the need for a better understanding of how and where it occurs. Without this understanding, finetuned models risk becoming less reliable, less robust, and ultimately less trustworthy, particularly in safety-critical applications where precision and predictability are paramount.
@@
  • We propose two approaches to analyze open-world forgetting, which are designed to assess semantic and appearance drift caused by the adaptation. We leverage the zero-shot classification capabilities of diffusion models to measure the semantic drift, and observe drastic performance drops (of over 60% for some classes). Appearance drift analysis confirms that customization leads to considerable changes in intra-class representation, color, and texture.
@@ 2 Related work
Text-to-image diffusion model adaptation.
Text-to-image (T2I) diffusion model adaptation is also referred to T2I personalization or subject-driven image generation. This aims to adapt a given model to a new concept by providing a few images and binding the new concept to a unique token. As a result, the adaptation model can generate various renditions of the new concept guided by text prompts. Depending on whether the adaptation method is finetuning the T2I model, they are categorized into two main streams. One of the most representative methods focuses on learning new concept tokens while freezing the T2I generative backbones. Textual Inversion (TI) (Gal et al., 2023) is a pioneering work focusing on finding new pseudo-words by performing personalization in the text embedding space. The following works (Dong et al., 2022; Daras & Dimakis, 2022; Voynov et al., 2023; Han et al., 2023a) continue to improve this technique stream. Another stream is finetuning the T2I generative models while updating the modifier tokens. One of the most representative methods is DreamBooth (Ruiz et al., 2023a), where the pre-trained T2I model learns to bind a modified unique identifier to a specific subject given 3$\sim$5 images, while it also updates the T2I model parameters. Custom Diffusion (Kumari et al., 2023) and other approaches (Han et al., 2023b; Chen et al., 2023b; Shi et al., 2023a) follow this pipeline and further improve the quality of the generation.
Finetuning methods often achieve state-of-the-art performance but introduce forgetting in large T2I models. While research focuses on improving new concept generation, it overlooks continuous model updating and forgetting mitigation. Recent works (Sun et al., 2024; Smith et al., 2023) address token forgetting but neglect other impacts of finetuning, such as semantic drifting in color, appearance, and visual recognition, which this paper explores.
@@
The main challenge of continual learning is to learn incrementally and accumulate knowledge of new data while preventing catastrophic forgetting, which is defined as a sudden drop in performance on previously acquired knowledge (McCloskey & Cohen, 1989; McClelland et al., 1995). The vast majority of studies on continual learning focus on, what we here call, closed-world forgetting, where the knowledge of the network can be represented by its performance on a limited set of classes (Lopez-Paz & Ranzato, 2017; De Lange et al., 2021; Masana et al., 2022). However, as argued in the introduction, the growing importance of starting from large pretrained models (also known as foundation models), which have a vast prior knowledge, requires new techniques to assess forgetting. The forgetting of large language models (LLMs) during continual finetuning has received some attention in recent years, showing the importance of pretraining to mitigate forgetting (Cossu et al., 2024), however, they mainly evaluate on down-stream-task performance Scialom et al. (2022). To the best of our knowledge, open-world forgetting has not yet been systematically analyzed for text-based image generation models which multi-modal nature can further worsen the impact of forgetting due to misalignment of the modalities.
@@ 3 Customization of Diffusion Models
In this section, we briefly introduce text-to-image (T2I) models and the two main customization methods we will evaluate during our analysis. In addition, we introduce an alternative regularization method to further mitigate forgetting.
@@ 3.1 Diffusion Models
Diffusion models are a class of generative models that generate data by gradually denoising a sample from a pure noise distribution. The process is modeled in two stages: a forward process and a reverse process. In the forward process, Gaussian noise is iteratively added to data samples, typically over $T$ steps, forming a Markov chain. At each step, the transition is defined as:
@@
where $\beta_{t}$ controls the noise schedule.
The reverse process denoises the data by learning the conditional probability $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, typically parameterized by a neural network $\epsilon_{\theta}(\mathbf{x}_{t},t)$ that predicts the added noise at each step. The model is trained by minimizing a simple mean-squared error between the true and predicted noise:
@@
Text-to-image diffusion models employ an additional conditioning vector $\mathbf{c}=\mathcal{E}(P)$ generated using a text encoder $\mathcal{E}$ and a text prompt $P$. These models have gained prominence for their ability to generate high-quality, diverse samples, often outperforming other generative models like GANs and VAEs in terms of mode coverage and sample quality (Ho et al., 2020; Song et al., 2021).
@@ 3.2 Customization approaches
Dreambooth (Ruiz et al., 2023b) enables personalization of diffusion models by finetuning them with a small set of images. It reuses an infrequent token of the vocabulary to represent a unique subject, allowing the model to generate images of the subject in varied contexts or styles. This approach induces language drift and reduced output diversity in the model, which is mitigated by replaying class-specific instances alongside the subject training, called prior preservation loss. The final training objective reads
@@
Custom diffusion (Kumari et al., 2023) is another approach aimed at efficiently finetuning diffusion models with minimal data and compute. This method observes that the cross-attention layer parameters undergo the most change during personalization, so they propose to only update the key and value projections in these layers. It introduces a token into the text-encoder representing a unique subject, rather than reusing an old one. By freezing the majority of the model’s parameters and focusing updates on a few key layers, Custom Diffusion facilitates rapid customization with less degradation in image quality. Prior preservation loss is maintained, since language drift is still experienced otherwise.
Customized Model Set In our experiments, we will evaluate Dreambooth and Custom Diffusion. We adapt both these models to ten different concepts based on 5 images per concept. The concepts are ‘lamp’, ‘vase’, ‘person2’,‘person3’,‘cat’,‘dog’,‘lighthouse’,‘waterfall’,‘bike’ and ‘car’ taken from CustomConcept101 (Kumari et al., 2023). We will refer to these ten models for both DreamBooth and Custom Diffusion as the Customized Model Set.
@@ 3.3 Drift Correction
The two studied approaches, Dreambooth (Ruiz et al., 2023b) and Custom Diffusion (Kumari et al., 2023), apply finetuning to adapt to the new data: they mainly focus on how good the learned model is on the target data, and do not study the possible detrimental effects for other classes. The Dreambooth method includes a method called prior regularization, which by replaying general instances of the concept being learned (see Eq. 3), helps to prevent the model from overfitting to the new data and ensures that the representation of the superclass remains stable. This same mitigation strategy is also applied in custom diffusion (Kumari et al., 2023).
@@ 4 Open-World Forgetting in Generative Model Adaptation
As an initial experiment, to assess open-world forgetting, we evaluate both the original unaltered model (called base model from now on) and the Customized Model Set on 10,000 user prompts from DiffusionDB (Wang et al., 2023) dataset (prompt examples are provided in Appendix B.1). Specifically, we measure the change of the resulting images using the cosine distance between CLIP-I encodings (Radford et al., 2021) when generating images with the same prompt and seed. Distances in the CLIP-I embedding are related to semantic similarity between images, with smaller distances indicating more similar visual content and larger distances suggesting more significant differences in the generated images. The distribution is plotted in Figure 2. For a detailed description of our experimental setup, please refer to Appendix B.1. It is important to note that a personalization method that does not alter the model would yield identical image outputs, resulting in a plot density concentrated at $1$.
When considering Figure 2a, we observe that even though most of the prompts from DiffusionDB are not related with the selected trained concepts, there is a significant part of the distribution that is shifted to the left. This shows indeed that the representations of the original model have changed. Furthermore, further analysis shows that open-world forgetting significantly alters the output in different ways, as illustrated by the samples in Figure 2a. For instance, a sampled pair from the most dissimilar outputs (purple triangle) shows a complete change in content, colors, and scene composition that no longer matches the prompt. In contrast, a very similar pair (yellow star) closely adheres to the original model’s output, with only changes in color or details. Interestingly, when looking at Figure 2b where we apply the proposed Drift Correction to DreamBooth, the distribution shifts to the right, showing that the drift has been reduced considerably.
Figure 2: Similarity (measured as cosine distance in CLIP-I embedding space) between models before and after adaptation. Each curve represents one of the 10 models from the Customized Model Set. a) Results with DreamBooth adaptation (includes prior regularization). b) Results with DreamBooth with Drift Correction. For more results see Appendix A.
@@ 4.1 Semantic Drift
Semantic drift refers to alterations in a model’s representation that cause the generation of semantically divergent content following customization. In the experiment depicted in Figure 2a, almost all prompts exhibit some level of drift, with a notable long tail of highly dissimilar generations. Many of these pronounced deviations have resulted in the generation of content that semantically no longer align with the input prompt.
To evaluate how semantic drift affects generative models, we use a straightforward approach: we utilize the model’s internal representations on different classification tasks (Mittal et al., 2023; Tang et al., 2023). It is based on a recent insight that showed that diffusion models can be directly applied for zero-class classification, by leveraging the conditional likelihood estimation of the model. Concretely, we use Diffusion Classifier (Li et al., 2023), where a posterior distribution over classes $\{\mathbf{c}_{i}\}_{i=1}^{N}$ is calculated as:
@@
Monte Carlo sampling is performed over $t_{i}$ and $\epsilon$ to obtain a classifier from a text-to-image model $\epsilon_{\theta}$.
While this method offers a simple, parameter-free approach to evaluating semantic drift, it is worth noting that alternative techniques have been proposed to assess the representation space of diffusion models. These include linear probing on activations (Xiang et al., 2023), analysis of hierarchical features (Mukhopadhyay et al., 2023), and methods requiring a preliminary likelihood maximization stage (Chen et al., 2023a). However, these alternatives often involve additional computational steps or are subject to specific settings, potentially limiting their applicability or introducing complexity to the evaluation process.
We conducted zero-shot classification experiments across multiple datasets spanning diverse domains to quantify the semantic drift of the models. We perform two measurements. First, we measure the average zero-shot classification score for the various models (the results are averaged over the 10 models of the Customized Model Set). Second, we establish the performance of the original pretrained model as the baseline, and measure the presence of semantic drift by calculating the drop in accuracy from the baseline. We also report the worst class drop which is the drop in accuracy of the class that has suffered the largest deterioration due to the adaptation. For further details on the classification method, please refer to Appendix B.2.
Table 1: Average zero-shot classification using the T2I models of the Customized Model Set for several image classification datasets. Worst class drop between parenthesis.
CIFAR10 STL10 Flowers Pets ObjectNet Food Aircraft
Base Model 81.60 93.00 50.00 86.87 28.50 71.09 23.40
@@
CD-DC 0.44 0.69 0.79
[…]
Furthermore, to measure generation quality, diversity, and alignment between image and text of the learned concepts, we employ metrics adapted from personalization methods: DINO, CLIP-I, and CLIP-T. In Table 2 we can see that the proposed regularization method DC does not negatively impact the image generation quality of the learned concepts. [^1] The results with standard deviations for Table 1 and 2 are provided in Appendix B.

### v2

@@ Abstract
Recent advances in diffusion models have significantly enhanced image generation capabilities. However, customizing these models with new classes often leads to unintended consequences that compromise their reliability. We introduce the concept of open-world forgetting to characterize the vast scope of these unintended alterations. Our work presents the first systematic investigation into open-world forgetting in diffusion models, focusing on semantic and appearance drift of representations. Using zero-shot classification, we demonstrate that even minor model adaptations can lead to significant semantic drift affecting areas far beyond newly introduced concepts, with accuracy drops of up to 60% on previously learned concepts. Our analysis of appearance drift reveals substantial changes in texture and color distributions of generated content. To address these issues, we propose a functional regularization strategy that effectively preserves original capabilities while accommodating new concepts. Through extensive experiments across multiple datasets and evaluation metrics, we demonstrate that our approach significantly reduces both semantic and appearance drift. Our study highlights the importance of considering open-world forgetting in future research on model customization and finetuning methods.
@@ 1 Introduction
Given the high-quality image generation capabilities of recent models (Shonenkov et al., 2023; Ramesh et al., 2022), a major focus of research has been on how to efficiently incorporate new content and adapt them to new tasks and domains. To address this challenge, state-of-the-art transfer learning methods have been introduced, including: finetuning approaches like DreamBooth (Ruiz et al., 2023) and CustomDiffusion (Kumari et al., 2023b) for learning new concepts, conditioning-based methods like ControlNet (Zhang & Agrawala, 2023) and IP-Adapters (Ye et al., 2023) for precise generation control, prompt-based techniques for semantic editing, and parameter-efficient methods like low-Rank Adaptations(LoRA) (Hu et al., 2022) for rapid adaptation with minimal overhead. These techniques enable models to learn new concepts effectively, even with few examples.
@@
Finetuning approaches for adapting diffusion models (Ruiz et al., 2023; Kumari et al., 2023b) mainly rely on transfer learning and primarily focus on finetuning model weights to accommodate newly introduced data. This weight update can have unintended consequences, altering the model’s existing knowledge and skills, or the alignment between language and visual content within the network, a phenomenon known as catastrophic forgetting in continual learning research (Kirkpatrick et al., 2017; De Lange et al., 2021). Traditionally, this phenomenon has been studied on what we term closed-world forgetting, where evaluation is limited to a fixed set of classes encountered in previously learned tasks or skills. This setting assumes a clear, predefined set of concepts to evaluate against. In contrast, modern foundation models introduce what we term open-world forgetting: degradation of the model’s capabilities across its vast, unconstrained knowledge space. Unlike closed-world settings, open-world forgetting is particularly challenging to measure since the model’s prior knowledge spans countless concepts, making it impossible to exhaustively evaluate what has been forgotten or altered during the adaptation process.
@@
In this paper, we focus on finetuning-based personalization methods for a case study of open-world forgetting. These techniques are especially interesting, as they only add little new knowledge to the network: a single new concept represented by a small set of typically 3-5 images. If open-world forgetting occurs even in this setting, where adaptation is limited to a handful of images, it strongly suggests that the phenomenon extends to larger-scale adaptations as well. Although one might expect that finetuning the model with such limited data would have minimal impact on the vast knowledge of the foundation model (e.g., Stable Diffusion), our analysis reveals that even these small updates can lead to highly detrimental consequences. As Figure 1 illustrates, finetuning can drastically alter the image representation of concepts seemingly unrelated to the training images. The complexity of the forgetting underscores the need for a better understanding of how and where it occurs. Without this understanding, finetuned models risk becoming less reliable, less robust, and ultimately less trustworthy, particularly in safety-critical applications where precision and predictability are paramount.
@@
  • We propose two approaches to analyze open-world forgetting, which are designed to assess semantic and appearance drift caused by the adaptation. We leverage the zero-shot classification capabilities of diffusion models to measure semantic drift, and observe drastic performance drops (of over 60% for some classes). Appearance drift analysis confirms that customization leads to considerable changes in intra-class representation, color, and texture.
@@ 2 Related work
Text-to-image diffusion model adaptation. Recent work has explored various approaches for personalizing text-to-image models. One line of research focuses on learning new concept tokens while freezing the generative backbone, pioneered by Textual Inversion Gal et al. (2023) and further developed in subsequent works Dong et al. (2022); Daras & Dimakis (2022); Voynov et al. (2023); Han et al. (2023a). Another approach involves finetuning the model while updating modifier tokens, exemplified by DreamBooth and Custom Diffusion. On these, the pre-trained T2I model learns to bind a modified unique identifier to a specific subject given 3$\sim$5 images, while it also updates the T2I model parameters.
Finetuning methods often achieve state-of-the-art performance but introduce forgetting in large T2I models, as we will demonstrate in this article. While research focuses on improving new concept generation, it overlooks continuous model updating and forgetting mitigation. Recent works (Sun et al., 2024; Smith et al., 2023) address token forgetting but neglect other impacts of finetuning, such as semantic drifting in color, appearance, and visual recognition, which this paper explores.
It should be noted that next to the training-time adaptation methods described above, recently also inference-time customization methods have been developed. These methods (Nair et al., 2023; Brack et al., 2023; Ruiz et al., 2024; Zhou et al., 2024) achieve adaptation by modifying the generation process without parameter updates, typically through diffusion guidance or prompt engineering. These methods are not considered in our paper since they do not consider parameter updates of the model.
@@
The vast majority of studies on Catastrophic forgetting focus on, what we here call, closed-world forgetting, where the knowledge of the network can be represented by its performance on a limited set of classes (Lopez-Paz & Ranzato, 2017; De Lange et al., 2021; Masana et al., 2022). However, as argued in the introduction, the growing importance of starting from large pretrained models (also known as foundation models), which have a vast prior knowledge, requires new techniques to assess forgetting. The forgetting of large language models (LLMs) during continual finetuning has received some attention in recent years, showing the importance of pretraining to mitigate forgetting (Cossu et al., 2024), however, they mainly evaluate on down-stream-task performance Scialom et al. (2022). To the best of our knowledge, open-world forgetting has not yet been systematically analyzed for text-based image generation models, whose multi-modal nature can further worsen the impact of forgetting due to misalignment of the modalities.
@@ 3 Customization of Diffusion Models
In this section, we briefly introduce T2I models and the two main customization methods we will evaluate during our analysis. In addition, we introduce an alternative regularization method to further mitigate forgetting.
@@ 3.1 Diffusion Models
Diffusion models generate data by gradually denoising samples from a noise distribution through a forward and reverse process. The forward process adds Gaussian noise over $T$ steps via a Markov chain:
@@
where $\beta_{t}$ controls the noise schedule. The reverse process learns to denoise by predicting the added noise through a neural network $\epsilon_{\theta}(x_{t},t)$, optimizing:
@@
Text-to-image diffusion models incorporate an additional conditioning vector $\mathbf{c}=\mathcal{E}(P)$ from a text encoder $\mathcal{E}$ and prompt $P$, enabling high-quality, diverse image generation conditioned on text descriptions.
@@ 3.2 Customization Approaches
Dreambooth (Ruiz et al., 2023) enables personalization of diffusion models by finetuning them with a small set of images. It reuses an infrequent token of the vocabulary to represent a unique subject, allowing the model to generate images of the subject in varied contexts or styles. This approach induces language drift and reduced output diversity in the model, which is mitigated by replaying class-specific instances alongside the subject training, called prior preservation loss. The final training objective reads
@@
Custom diffusion (Kumari et al., 2023b) follows a similar approach but updates only the key and value projections in cross-attention layers, introducing a new token rather than reusing existing ones. Prior preservation loss is maintained, since language drift is still experienced otherwise.
Customized Model Set In our experiments, we will evaluate Dreambooth and Custom Diffusion. We adapt both these models to 10 different concepts based on 5 images per concept. The concepts are ‘lamp’, ‘vase’, ‘person2’,‘person3’,‘cat’,‘dog’,‘lighthouse’,‘waterfall’,‘bike’ and ‘car’ taken from CustomConcept101 (Kumari et al., 2023b). We will refer to these ten models for both DreamBooth and Custom Diffusion as the Customized Model Set. We have experimentally found that 10 concepts is enough for robust generalization. For a scale analysis, please see Appendix A.2.
@@ 3.3 Drift Correction
The two studied approaches, Dreambooth and Custom Diffusion, apply finetuning to adapt to the new data: they mainly focus on how good the learned model is on the target data, and do not study the possible detrimental effects for other classes. The Dreambooth method includes a method called prior regularization, which by replaying general instances of the concept being learned (see Eq. 3), helps to prevent the model from overfitting to the new data and ensures that the representation of the superclass remains stable. This same mitigation strategy is also applied in custom diffusion (Kumari et al., 2023b).
@@ 4 Open-World Forgetting in Generative Model Adaptation
As an initial experiment, to assess open-world forgetting, we evaluate both the original unaltered model (called base model from now on) and the Customized Model Set on 10,000 user prompts from DiffusionDB (Wang et al., 2023) dataset (prompt examples and further experimental details are provided in Appendix E.1). Specifically, in Figure 2, we measure the change of the resulting images using the cosine distance between CLIP-I encodings (Radford et al., 2021) when generating images with the same prompt and seed. Distances in the CLIP-I embedding are related to semantic similarity between images, with smaller distances indicating more similar visual content and larger distances suggesting more significant differences in the generated images. A personalization method that does not alter the model would yield identical image outputs, resulting in a plot density concentrated at $1$.
When considering Figure 2a, we observe that even though most of the prompts from DiffusionDB are not related with the selected trained concepts, there is a significant part of the distribution that is shifted to the left. This shows indeed that the representations of the original model have changed. Furthermore, further analysis shows that open-world forgetting significantly alters the output in different ways, as illustrated by the samples in Figure 2a. For instance, a sampled pair from the most dissimilar outputs (purple triangle) shows a complete change in content, colors, and scene composition that no longer matches the prompt. In contrast, a very similar pair (yellow star) closely adheres to the original model’s output, with only changes in color or details [^1] Similar curves are observed for different model families (see Appendix A.3).. Interestingly, when looking at Figure 2b where we apply the proposed Drift Correction to DreamBooth, the distribution shifts to the right, showing that the drift has been reduced considerably.
Figure 2: Similarity (measured as cosine distance in CLIP-I embedding space) between models before and after adaptation. Each curve represents one of the 10 models from the Customized Model Set. a) Results with DreamBooth adaptation (includes prior regularization). b) Results with DreamBooth with Drift Correction. For more results see Appendix F.
@@ 4.1 Semantic Drift
Semantic drift represents fundamental changes in how the model understands and represents concepts, causing the generation of semantically divergent content following customization. In the experiment depicted in Figure 2a, almost all prompts exhibit some level of drift, with a notable long tail of highly dissimilar generations. Many of these pronounced deviations have resulted in the generation of content that semantically no longer align with the input prompt.
To evaluate how semantic drift affects generative models, we use a straightforward approach: we utilize the model’s internal representations on different classification tasks (Mittal et al., 2023; Tang et al., 2023). It is based on a recent insight that diffusion models can be directly applied for zero-class classification by leveraging the conditional likelihood estimation of the model. This provides a direct measure of how well the model maintains its understanding of various concepts. Concretely, we use Diffusion Classifier (Li et al., 2023), where a posterior distribution over classes $\{\mathbf{c}_{i}\}_{i=1}^{N}$ is calculated as:
@@
Monte Carlo sampling is performed over $t_{i}$ and $\epsilon$ to obtain a classifier from the model $\epsilon_{\theta}$.
While this method offers a simple, parameter-free approach to evaluating semantic drift, it is worth noting that alternative techniques have been proposed to assess the representation space of diffusion models. These include linear probing on activations (Xiang et al., 2023), analysis of hierarchical features (Mukhopadhyay et al., 2023), and methods requiring a preliminary likelihood maximization stage (Chen et al., 2023). However, these alternatives often involve additional computational steps or are subject to specific settings, potentially limiting their applicability or introducing complexity to the evaluation process.
We conducted zero-shot classification experiments across multiple datasets spanning diverse domains to quantify the semantic drift of the models. We perform two measurements. First, we measure the average zero-shot classification score for the various models (the results are averaged over the 10 models of the Customized Model Set). Second, we establish the performance of the original pretrained model as the baseline, and measure the presence of semantic drift by calculating the drop in accuracy from the baseline. We also report the worst class drop which is the drop in accuracy of the class that has suffered the largest deterioration due to the adaptation. For further details on the classification method, please refer to Appendix E.2.
Table 1: Average zero-shot classification using the T2I models of the Customized Model Set for several image classification datasets, with concept and prompt fidelity metrics. Worst class drop between parenthesis. Concept fidelity (DINO, CLIP-I) and prompt fidelity (CLIP-T). Drift Correction maintains fidelity across metrics.
CIFAR10 STL10 Flowers Pets ObjectNet Food Aircraft DINO CLIP-I CLIP-T
Base Model 81.60 93.00 50.00 86.87 28.50 71.09 23.40 - - -
@@
Furthermore, to measure generation quality, diversity, and alignment between image and text of the learned concepts, we employ metrics adapted from personalization methods: DINO, CLIP-I, and CLIP-T. In Table 1 we can see that the proposed regularization method DC does not negatively impact the image generation quality of the learned concepts.
Standard deviations and more zero-shot evaluation for more recent methods, including BOFT Liu et al. (2024) and SVDiff Han et al. (2023b) are provided in Appendix B.1.